# Diffusion-VLA: Scaling Robot Foundation Models via Unified Diffusion and Autoregression

### v1

@@ Abstract
In this paper, we present DiVLA, a novel framework that seamlessly combines the autoregression model with the diffusion model for learning visuomotor policy. Central to our approach is a next-token prediction objective, enabling the model to reason effectively over the user’s query in the context of current observations. Subsequently, a diffusion model is attached to generate robust action outputs. To enhance policy learning through self-reasoning, we introduce a novel reasoning injection module that integrates reasoning phrases directly into the policy learning process. The whole framework is simple and flexible, making it easy to deploy and upgrade.
We conduct extensive experiments using multiple real robots to validate the effectiveness of DiVLA. Our tests include a challenging factory sorting task, where DiVLA successfully categorizes objects, including those not seen during training. We observe that the reasoning module enhances interpretability, allowing observers to understand the model’s thought process and identify potential causes of policy failures. Additionally, we test DiVLA on a zero-shot bin-picking task, achieving 63.7% accuracy on 102 previously unseen objects. Our method demonstrates robustness to visual changes, such as distractors and new backgrounds, and easily adapts to new embodiments. Furthermore, DiVLA can follow novel instructions and retain conversational ability. Notably, DiVLA is data-efficient and fast at inference; our smallest DiVLA-2B runs 82Hz on a single A6000 GPU and can train from scratch on less than 50 demonstrations for a complex task. Finally, we scale the model from 2B to 72B parameters, showcasing improved generalization capabilities with increased model size.
@@ 1 Introduction
Autoregressive models are predominant in Large Language Models, operating by sequentially predicting discrete tokens, where each token generation depends on the preceding ones. This approach has also influenced robotics, leading to the development of Vision-Language-Action (VLA) models, which frame action prediction as a next-token prediction task. While these models, i.e., RT-2 [9] and OpenVLA [26], have achieved notable success, they face intrinsic limitations. First, discretizing continuous action data into fixed-size tokens disrupts the coherence and precision of actions. Second, NTP is inherently inefficient for action generation, especially in real-time robotic applications where performance is critical.
@@
Meanwhile, building on the success of diffusion models in content generation [43, 36, 38, 15, 16], diffusion-based models for learning visuomotor policies [10] have gained significant popularity over the past two years. Numerous methods have demonstrated strong performance in manipulation tasks by modeling action sequence generation as a noise-denoising process. This approach better captures the multimodal nature of robotic actions and enables faster sequence generation compared to NTP-based VLA models. However, despite the advantages of diffusion models for policy learning, they lack the reasoning capabilities crucial for VLA models to solve complex tasks effectively, a component that evidently improves LLMs.
This motivates us to raise the question: can we bring together the advantages of both parties, specifically the reasoning power of autoregressive models and the robustness of high-frequency action generation offered by diffusion models?
@@
In this work, we propose a unified model, named DiffusionVLA (DiVLA in short), that integrates autoregression with a diffusion model. The autoregressive component provides reasoning over the query, while the diffusion model controls the robot. Specifically, DiVLA builds upon a pre-trained Vision-Language Model (VLM), retaining its autoregressive capabilities for text-based reasoning. We extend this foundation by integrating a diffusion model that facilitates learning robotic actions through a noise-denoising process. This setup empowers DiVLA to achieve both language-driven reasoning and robust action generation in robotic contexts. However, simply combining these elements does not fully exploit the reasoning potential, as there is often an implicit gap between logical reasoning and actionable robot policies. To bridge this gap, we propose a reasoning injection module, which reuses reasoning outputs and embeds them directly into the policy head, thus enriching the policy learning process with explicit reasoning signals. This innovation allows us to directly incorporate reasoning into action generation, enhancing the model’s dexterity, robustness, and generalization across various scenarios. Our experiments confirm that DiVLA delivers the following advantages:
  • Fast Inference Speed: DiVLA-2B achieves an inference rate of 82Hz, and DiVLA-7B operates at 42Hz on a single A6000 GPU, ensuring real-time responsiveness in high-demand environments.
  • Enhanced Visual Generalization: DiVLA is unaffected by visual distractions or novel backgrounds, demonstrating robustness in visually dynamic settings.
  • Generalizable Reasoning Capability: DiVLA can recognize and categorize previously unseen objects accurately, showcasing its ability to generalize reasoning across novel inputs.
@@
  • Adaptability to Novel Instructions and Conversational Capability: Our approach can interpret and execute complex, novel instructions while maintaining conversational fluency, offering a versatile response range in interactive scenarios.
@@
  • Generalization to Other Embodiment: DiVLA readily fine-tunes for deployment on real bimanual robots, achieving high performance with minimal adjustment, proving its adaptability across diverse robotic embodiments.
@@
  • Scalability: We provide a scalable family of models — DiVLA-2B, 7B, and 72B — demonstrating that generalization and performance improve with model size, adhering to established scaling laws.
We believe that DiVLA offers a novel perspective on the learning and generalizing robot policy.
@@ 2 Related Works
Autoregression models. Predicting the next token has been regarded as a key approach toward general artificial intelligence due to its success in training language models [48, 49, 1, 4, 47]. RT-2 [9] pioneered the application of next-token prediction in robot learning which predicts actions by converting continuous actions into discrete tokens for learning robotic motion. Building on this, OpenVLA [26] introduced an open-source, improved, and smaller version of RT-2 [9] with a similar approach, while ECoT [64] developed a chain-of-thought method. However, research has shown that next-token prediction may not be the optimal approach for robot models, especially when adapting to various embodiments. In this work, we leverage the strength of next-token prediction specifically for reasoning tasks.
@@
Diffusion models. Diffusion models have become dominant in the field of visual generation. The Diffusion Policy [10] extends the application of diffusion models to robot learning, demonstrating their effectiveness in handling multimodal action distributions. Subsequent work has advanced the Diffusion Policy [71, 55, 39, 42, 50, 51, 5, 6, 11, 30] by applying it to 3D settings [24, 68, 66, 62], scaling it up [74], increasing its efficiency [21, 56], and introducing architectural innovations. For instance, TinyVLA [57] integrates diffusion models with lightweight vision-language models, while $pi_{0}$ [7] leverages flow matching rather than diffusion for action generation. Our approach introduces reasoning—a crucial element in language models—into the diffusion-based vision-language-action (VLA) model.
@@
Robot foundation models. Existing works [3, 8, 9, 14, 22, 20, 12, 75] leverage RL [63, 41, 19, 32, 45, 67, 65] and LLM [29, 58, 73, 2, 40, 18, 23] to decouple the multimodal understanding and embodied control. Another line of research leverages pre-trained Vision-Language Models (VLMs) and fine-tunes them directly on robotic data [9, 64, 26, 57]. Our work follows this approach, unifying autoregressive and diffusion models for both reasoning and manipulation tasks.
@@
Unified auto-regressive model and image generation. Recent work has focused on unifying multimodal understanding with image generation. These efforts include using auto-regressive methods to generate images [28, 46, 54, 52, 33], diffusion models to produce text, or combining both approaches into unified models [13, 31, 70, 59], such as Show-O [61], Transfusion [72], and Vila-U [60]. Unlike these methods, our framework explores unifying next-token prediction with diffusion models, which enhances the robot model’s reasoning abilities. This reasoning capability, in turn, improves the model’s generalization to various tasks and environments.
@@ 3 Methodology
3.1 Architecture
Given any sequence of interleaved images, text, and video, we first encode the images into dense visual features using SigLIP [69]. These encodings are then transformed into a fixed number of $N$ visual embeddings through a Transformer. It’s worth noting that typical visual inputs in robot learning often include multiple camera views. To manage this, we applied the shared SigLIP visual backbone to each view and subsequently concatenated the resulting visual tokens.
For vision-language processing, we leveraged Qwen2-VL [53], a state-of-the-art vision-language model available in three sizes: 2B, 7B, and 72B parameters. We initialized the VLM backbone with the publicly released checkpoint. It is also possible to use any other pre-trained VLM as backbone, since we decouple the vision-language understanding with action generation, making the overall architecture flexible to fit for advanced new models.
Action decoder. We employ a latent diffusion model to decode visual and instruction embeddings into actions. Specifically, we generate a set of tokens using an LLM and feed these tokens into the diffusion model as conditioning inputs for action decoding. Our architecture follows the design of the standard Diffusion Policy [10], with randomly initialized model weights. This component also incorporates reasoning from the LLM, which we describe in detail below. An MLP layer is attached to the last layer at the bottom of the action decoder to predict the robot’s joint space. If multiple embodiments are evolved, instead of making a copy of a separate action decoder [34], we simply initialized a new MLP layer for training and evaluation. This step ensures that the knowledge from the pre-trained data is preserved and thus can quickly adapt to a new embodiment.
Our reasoning injection module operates by taking the final embedding from the tokenized output of the reasoning component and directly injecting it into the policy model through Feature-wise Linear Modulation (FiLM) [37]. This injection technique, inspired by methods in RT-1 [8] and YAY [44], enables us to modulate the policy network’s layers based on the reasoning signal. We refer to this process as “injection” because, in our design, the policy network focuses primarily on action-specific tokens, while the reasoning module functions as an auxiliary enhancement, providing contextual depth without dominating the primary decision-making flow. This approach ensures that reasoning is not only present but actively utilized during policy model training.
3.2 Model Design Choices
View-adaptive tokenization. As experiments in robot policy learning commonly use multiple camera views. As the number of views increases, the count of visual tokens also rises, leading to significant computational overhead due to the complexity of self-attention. To address this, we propose a view-adaptive tokenization method that reduces the number of visual tokens, optimizing computational efficiency.
Our approach is straightforward yet effective: we significantly reduce the number of tokens for wrist camera views to just 16, which is over 20 times fewer than those for extrinsic views. This reduction is motivated by the distinct role of the wrist camera, which primarily assists in precise localization for grasping tasks. High-resolution processing of wrist views, while detailed, is often inefficient and consumes resources that could be better allocated to processing more critical visual signals from external perspectives.
This optimization is particularly beneficial in dual-arm robotic systems, where two wrist cameras are active simultaneously. By reducing the computational load associated with wrist views, we free up resources to handle other essential tasks, enabling smoother, more efficient operation and potentially enhancing the overall responsiveness and precision of the robotic system.
Training objectives. Given a batch of input sequences, the overall training loss is formulated as a combination of the diffusion loss and the next-token prediction loss: $L=L_{diff}+\alpha L_{ntp}$, where $\alpha$ is the hyper-parameters weighting the loss term $L_{ntp}$. In our observations, the magnitude of $L_{ntp}$ consistently remains about ten times smaller than that of $L_{diff}$. To balance the contribution of each component to the overall loss, we typically set $\alpha=1$ in all experiments. This adjustment ensures that both terms are comparably weighted in the training process, allowing the model to learn effectively from both action and next-token prediction tasks.
Pretraining Data. We consider OXE [35] and Droid [25] dataset for pretraining. We use Droid data to pre-train DiVLA-2B and DiVLA-7B. For DiVLA-72B, we use OXE and Droid together for pre-train. The original Droid data contains only robotic actions, paired partially with observations and language instructions. These data contain only robotic actions, paired partially with observations and language instructions. To enhance our model’s ability to generalize with language, we leverage GPT-4o to automatically transform these data into a form that includes reasoning. Consequently, the network architecture remains consistent across both the pre-training and fine-tuning stages.
Figure 1: Environmental setup for the Franka Robot and experimental configuration for sorting/bin-picking. Left: For factor sorting tasks, (a) The target sorting box is divided into four distinct sectors, each designated for one of the following categories: stuffed toys, hex keys, knit gloves, and toy cars, (c) The seen objects in the train data, (d) mixing the seen and unseen object for evaluation, (e) cluttered scene for seen objects, (f) cluttered scene for mixing seen and unseen objects. Middle: We use a Franka robot arm equipped with two external Zed cameras and a Realsense 435i wrist camera. Right: The setup for zero-shot bin picking.
4 Experiments
In this section, we examine the effectiveness of DiVLA for embodied control. In Section 4.2, we compare DiVLA  against other state-of-the-art models within a standard multi-task setting, assessing its performance in both in-distribution and out-of-distribution scenarios. In Section 4.3, we evaluate DiVLA  in the challenging sorting task, showcasing its remarkable performance and illustrating how reasoning enables the model to analyze robot actions and sort items accurately. In Section 4.4, we showcase DiVLA’s impressive generalization abilities in a zero-shot bin-picking task involving over 102 unseen objects. In Section 4.5, we illustrate DiVLA’s adaptability to new embodiments, successfully completing complex tasks that require bimanual coordination. We provide experimental setup and implementation details in the Appendix.
4.1 Experimental Setup.
Implementation details and pretraiend data. The model is pre-trained on the Droid [25] dataset. We then finetune our model on evaluation tasks, similar to the setting as $\pi_{0}$ [7]. We use LoRA [17] to fine-tune the VLM models. We use 2e-5 as a fixed learning rate to pretrain the model.
Data for finetuning. We explore four experimental settings: sorting, bin picking, multi-task learning, and table bussing. The first three settings are conducted with the Franka robot, while the table bussing task utilizes the bimanual AgileX robot. Our dataset includes 500 trajectories for the sorting task and 580 trajectories for multi-task learning. The bin picking task is designed as a zero-shot task, so no training data was collected for it. For the table bussing task, we gathered 400 trajectories, where objects are randomly placed on the table, often overlapping with each other. During the fine-tuning stage, all data corresponding to the same embodiment are trained together. For instance, the sorting and multi-task data are combined for training purposes.
Figure 2: Experimental results for sorting. We compared our DiVLA with Diffusion Policy, Octo, TinyVLA, and OpenVLA. DiVLA achieves the highest average success rate, outperforming the runner-up OpenVLA by 20.9%.
@@ 4 Experiments
Diffusion Policy [10] - 66.7 36.4 0 36.4 0 27.9 11.1 11.1 0 22.2 0 8.9
ScalingDP-1B [74] - 66.7 36.4 27.3 37.0 18.2 35.2 33.3 33.3 0 22.2 22.2 22.2
TinyVLA [57] - 72.7 45.5 36.4 45.5 27.3 45.5 44.4 44.4 11.1 22.2 22.2 28.9
Octo [34] 970K 57.6 27.3 9.1 0 27.3 24.3 44.4 11.1 11.1 0 22.2 17.8
OpenVLA-7B [26] 970K 69.7 18.2 18.2 36.4 54.5 39.4 55.6 11.1 0 33.3 33.3 26.7
@@
DiVLA-7B 39K 100 100 72.7 81.8 100 90.9 66.7 88.9 55.6 77.8 66.7 71.0

### v2

# Diffusion-VLA: Generalizable and Interpretable Robot Foundation Model via Self-Generated Reasoning

@@ Abstract
In this paper, we present DiffusionVLA, a novel framework that integrates autoregressive reasoning with diffusion policies to address the limitations of existing methods: while autoregressive Vision-Language-Action (VLA) models lack precise and robust action generation, diffusion-based policies inherently lack reasoning capabilities. Central to our approach is autoregressive reasoning — a task decomposition and explanation process enabled by a pre-trained VLM — to guide diffusion-based action policies. To tightly couple reasoning with action generation, we introduce a reasoning injection module that directly embeds self-generated reasoning phrases into the policy learning process. The framework is simple, flexible, and efficient, enabling seamless deployment across diverse robotic platforms.
We conduct extensive experiments using multiple real robots to validate the effectiveness of DiVLA. Our tests include a challenging factory sorting task, where DiVLA successfully categorizes objects, including those not seen during training. The reasoning injection module enhances interpretability, enabling explicit failure diagnosis by visualizing the model’s decision process. Additionally, we test DiVLA on a zero-shot bin-picking task, achieving 63.7% accuracy on 102 previously unseen objects. Our method demonstrates robustness to visual changes, such as distractors and new backgrounds, and easily adapts to new embodiments. Furthermore, DiVLA can follow novel instructions and retain conversational ability. Notably, DiVLA is data-efficient and fast at inference; our smallest DiVLA-2B runs 82Hz on a single A6000 GPU. Finally, we scale the model from 2B to 72B parameters, showcasing improved generalization capabilities with increased model size.
Machine Learning, ICML
Figure 1: Our proposed DiffusionVLA model unifies autoregressive and diffusion modeling to enable self-reasoning and robot policy learning. This approach generalizes effectively to visual changes, supports zero-shot bin picking, adapts to new robot morphologies, performs visual question-answering, and generates actions with high speed.
@@ 1 Introduction
Recently, Vision-Language-Action (VLA) models have emerged as a promising direction in robotics (Bu et al., 2024; Liu et al., 2024; Li et al., 2024c, a; Zhang et al., 2024; Zheng et al., 2025; Pertsch et al., 2025; Brohan et al., 2023; Kim et al., ; Wen et al., 2024; Black et al., 2024; Octo Model Team et al., 2024). Among these VLAs, a prominent approach frames action prediction as a next-token prediction (NTP) task, mirroring the dominant autoregressive paradigm in Large Language Models (LLMs), which operates by sequentially predicting discrete tokens, with each token’s generation conditioned on the preceding ones. While these models, such as RT-2 (Brohan et al., 2023) and OpenVLA (Kim et al., ), have demonstrated notable success, they suffer from inherent limitations. First, discretizing continuous action data into fixed-size tokens can disrupt the coherence and precision of actions. Second, NTP is inherently inefficient for action generation, particularly in real-time robotic applications where performance is critical.
@@
Meanwhile, building on the success of diffusion models in content generation (Rombach et al., 2022; Peebles & Xie, 2023; Podell et al., 2023; Ho et al., 2020, 2022), diffusion-based models for learning visuomotor policies (Chi et al., 2023) have gained significant popularity over the past two years. Numerous methods have demonstrated strong performance in manipulation tasks by modeling action sequence generation as a noise-denoising process. This approach better captures the multimodal nature of robotic actions and enables faster sequence generation compared to NTP-based VLA models. However, despite the advantages of diffusion models for policy learning, they lack the reasoning capabilities crucial for VLA models to solve complex tasks effectively, a component that evidently improves LLMs. This motivates us to raise the question: can we bring together the advantages of both parties, specifically the reasoning power of autoregressive models and the robustness of high-frequency action generation offered by diffusion models?
@@
In this work, we propose a unified model, named DiffusionVLA (DiVLA in short), that integrates autoregression with a diffusion model. The autoregressive component provides reasoning over the query, while the diffusion model controls the robot. Specifically, DiVLA builds upon a pre-trained Vision-Language Model (VLM), retaining its autoregressive capabilities for text-based reasoning. We extend this foundation by integrating a diffusion model that facilitates learning robotic actions through a noise-denoising process. This setup empowers DiVLA to achieve both language-driven reasoning and robust action generation in robotic contexts. However, simply combining these elements does not fully exploit the reasoning potential, as there is often an implicit gap between logical reasoning and actionable robot policies. To bridge this gap, we propose a reasoning injection module, which reuses reasoning outputs and embeds them directly into the policy head, thus enriching the policy learning process with explicit reasoning signals. This innovation allows us to directly incorporate reasoning into action generation, enhancing the model’s dexterity, robustness, and generalization across various scenarios. Our experiments confirm that DiVLA delivers the following advantages: 1) Visual generalization via self-generated reasoning: DiVLA can recognize and categorize previously unseen objects via self-generated reasoning, showcasing its ability to generalize to novel visual inputs.
2) Strong action interpretability: Our reasoning injection module provides insights into the end-to-end policy’s decision-making, explaining robot actions and facilitating failure analysis.
@@
3) Adaptability to novel instructions and conversational capability: Our approach can execute novel instructions while maintaining conversational fluency, offering a versatile response range in interactive scenarios.
@@
4) Fast adaptation to other embodiment: DiVLA can quickly fine-tune for deployment on new embodiment, such as bimanual robots, achieving high performance with minimal adjustments.
5) Fast inference speed: With inference rates of 82Hz for DiVLA-2B and 42Hz for DiVLA-7B on a single A6000 GPU, our method ensures real-time responsiveness even in high-demand environments.
6) Enhanced visual generalization: DiVLA is unaffected by visual distractions or novel backgrounds, demonstrating robustness in visually dynamic settings.
@@
7) Scalability: The scalable DiVLA family (2B, 7B, and 72B) demonstrates that generalization and performance improve with model size, consistent with established scaling laws.
This work introduces a novel end-to-end framework for robot policy learning. By unifying an autoregressive objective with a diffusion model, our approach enables the model to both “talk” and “act.” Critically, our proposed reasoning injection module promotes generalization to novel objects and provides interpretable actions. While the concept is simple and straightforward, the combination of ingredients and the application to robot learning is novel. Empirical evaluation across multiple complex real-world robot tasks demonstrates a level of generality surpassing that of previous single-model approaches.
@@ 2 Related Works
Autoregression models. Predicting the next token has been regarded as a key approach toward general artificial intelligence due to its success in training language models (Touvron et al., 2023a, b; Achiam et al., 2023; Bi et al., 2024; Team, 2024). RT-2 (Brohan et al., 2023) pioneered the application of next-token prediction in robot learning which predicts actions by converting continuous actions into discrete tokens for learning robotic motion. Building on this, OpenVLA (Kim et al., ) introduced an open-source, improved, and smaller version of RT-2 (Brohan et al., 2023) with a similar approach, while ECoT (Zawalski et al., 2024) developed a chain-of-thought method. However, research has shown that next-token prediction may not be the optimal approach for robot models, especially when adapting to various embodiments. In this work, we leverage the strength of next-token prediction specifically for reasoning tasks.
@@
Diffusion models. Diffusion models have become dominant in the field of visual generation. The Diffusion Policy (Chi et al., 2023) extends the application of diffusion models to robot learning, demonstrating their effectiveness in handling multimodal action distributions. Subsequent work has advanced the Diffusion Policy (Zhao et al., ; Wang et al., 2024d; Prasad et al., 2024; Reuss et al., 2024; Uehara et al., 2024a, b; Black et al., 2023a, b; Dasari et al., 2024; Lin et al., 2024) by applying it to 3D settings (Ke et al., 2024; Ze et al., 2024b, a; Yan et al., 2024), scaling it up (Zhu et al., 2024), increasing its efficiency (Jia et al., ; Wang et al., 2024e), and introducing architectural innovations. For instance, TinyVLA (Wen et al., 2024) integrates diffusion models with lightweight vision-language models, while $\pi_{0}$ (Black et al., 2024) leverages flow matching rather than diffusion for action generation. Our approach introduces reasoning—a crucial element in language models—into the diffusion-based vision-language-action (VLA) model.
@@
Robot foundation models. Existing works (Belkhale et al., 2024; Brohan et al., 2022, 2023; Gu et al., 2023; Jiang et al., 2022; Jang et al., 2022; Fu et al., 2024; Huang et al., ) leverage RL (Yuan et al., 2024) and LLM (Liang et al., 2023) to decouple the multimodal understanding and embodied control. Another line of research leverages pre-trained Vision-Language Models (VLMs) and fine-tunes them directly on robotic data (Brohan et al., 2023; Zawalski et al., 2024; Kim et al., ; Wen et al., 2024). Our work follows this approach, unifying autoregressive and diffusion models for both reasoning and manipulation tasks.
@@
Unified auto-regressive model and image generation. Recent work has focused on unifying multimodal understanding with image generation. These efforts include using auto-regressive methods to generate images (Li et al., 2024b; Sun et al., 2024; Wang et al., 2024c, a), diffusion models to produce text, or combining both approaches into unified models (Ge et al., 2024; Lu et al., 2024; Zhao et al., 2024; Wu et al., 2024a), such as Show-O (Xie et al., 2024), Transfusion (Zhou et al., 2024), and Vila-U (Wu et al., 2024b). Unlike these methods, our framework explores unifying next-token prediction with diffusion models, which enhances the robot model’s reasoning abilities. This reasoning capability, in turn, improves the model’s generalization to various tasks and environments.
Figure 2: Environmental Setup for the Franka Robot and Experimental Configuration for Factory Sorting. Left: For factor sorting tasks, (a) The target sorting box is divided into four distinct sectors, each designated for one of the following categories: stuffed toys, hex keys, knit gloves, and toy cars, (c) The seen objects in the train data, (d) mixing the seen and unseen object for evaluation, (e) cluttered scene for seen objects, (f) cluttered scene for mixing seen and unseen objects. Middle: We use a Franka robot arm equipped with two external Zed cameras and a Realsense 435i wrist camera. Right: The setup for zero-shot bin picking.
@@ 3 Methodology
Diffusion Policy (Chi et al., 2023) - 66.7 36.4 0 36.4 0 27.9 11.1 11.1 0 22.2 0 8.9
TinyVLA (Wen et al., 2024) - 72.7 45.5 36.4 45.5 27.3 45.5 44.4 44.4 11.1 22.2 22.2 28.9
Octo (Octo Model Team et al., 2024) 970K 57.6 27.3 9.1 0 27.3 24.3 44.4 11.1 11.1 0 22.2 17.8
OpenVLA-7B (Kim et al., ) 970K 69.7 18.2 18.2 36.4 54.5 39.4 55.6 11.1 0 33.3 33.3 26.7
@@
Figure 3: Experimental Results for Factory Sorting. We compared our DiVLA with Diffusion Policy, Octo, TinyVLA, and OpenVLA. DiVLA achieves the highest average success rate, outperforming the runner-up OpenVLA by 20.9%.
3.1 Architecture
Given any sequence of interleaved images, text, and video, we first encode the images into dense visual features using SigLIP (Zhai et al., 2023). These encodings are then transformed into a fixed number of $N$ visual embeddings through a Transformer. It’s worth noting that typical visual inputs in robot learning often include multiple camera views. To manage this, we applied the shared SigLIP visual backbone to each view and subsequently concatenated the resulting visual tokens.
For vision-language processing, we leveraged Qwen2-VL (Wang et al., 2024b), a state-of-the-art vision-language model available in three sizes: 2B, 8B, and 72B parameters. We initialized the VLM backbone with the publicly released checkpoint. It is also possible to use any other pre-trained VLM as backbone, since we decouple the vision-language understanding with action generation, making the overall architecture flexible to fit for advanced new models.
Projection layer for action tokens. Following the final embedding layer of the VLM, a fixed number of action tokens are generated. These tokens are then fed into a projection module, comprised of two MLP layers with LayerNorm. This projection module functions similarly to those found in conventional vision-language models like LLaVA (Liu et al., 2023b, a), bridging the VLM’s output embedding to the diffusion model and aligning their output dimensions. The diffusion model itself adheres to the standard Diffusion Policy design (Chi et al., 2023), with randomly initialized weights This component also incorporates reasoning from the LLM, which we describe in detail below. An MLP layer is attached to the last layer at the bottom of the action decoder to predict the robot’s joint space. If multiple embodiments are evolved, instead of making a copy of a separate action decoder (Octo Model Team et al., 2024), we simply initialized a new MLP layer for training and evaluation. This step ensures that the knowledge from the pre-trained data is preserved and thus can quickly adapt to a new embodiment.
Our reasoning injection module operates by taking the final embedding from the tokenized output of the reasoning component and directly injecting it into the policy model through Feature-wise Linear Modulation (FiLM) (Perez et al., 2018). This injection technique, inspired by methods in RT-1 (Brohan et al., 2022) and YAY (Shi et al., 2024), enables us to modulate the policy network’s layers based on the reasoning signal. We refer to this process as “injection” because, in our design, the policy network focuses primarily on action-specific tokens, while the reasoning module functions as an auxiliary enhancement, providing contextual depth without dominating the primary decision-making flow. This approach ensures that reasoning is not only present but actively utilized during policy model training.
3.2 Model Design Choices
Training objectives. Given a batch of input sequences, the overall training loss is formulated as a combination of the diffusion loss and the next-token prediction (ntp) loss: $L=L_{diff}+\alpha L_{ntp}$, where $\alpha$ is the hyper-parameters weighting the loss term $L_{ntp}$. In our observations, the magnitude of $L_{ntp}$ consistently remains about ten times smaller than that of $L_{diff}$. To balance the contribution of each component to the overall loss, we typically set $\alpha=10$ in all experiments. This adjustment ensures that both terms are comparably weighted in the training process, allowing the model to learn effectively from both action and next-token prediction tasks.
Pretraining Data. We consider OXE (O’Neill et al., 2023) and Droid (Khazatsky et al., 2024) dataset for pretraining. We use Droid data to pre-train DiVLA-2B and DiVLA-7B. Because larger models typically needs more data for training, we use OXE and Droid together for pre-training DiVLA-72B. The original Droid data contains only robotic actions, paired partially with observations and language instructions. These data contain only robotic actions, paired partially with observations and language instructions. To enhance our model’s ability to generalize with language, we leverage GPT-4o to automatically transform these data into a form that includes reasoning. Consequently, the network architecture remains consistent across both the pre-training and fine-tuning stages.
4 Experiments
In this section, we examine the effectiveness of DiVLA for embodied control. In Section 4.2, we compare DiVLA  against other state-of-the-art models within a standard multi-task setting, assessing its performance in both in-distribution and out-of-distribution scenarios. In Section 4.3, we evaluate DiVLA  in the challenging factory sorting task, showcasing its remarkable performance and illustrating how reasoning enables the model to analyze robot actions and sort items accurately. In Section 4.5, we showcase DiVLA’s impressive generalization abilities in a zero-shot bin-picking task involving over 102 unseen objects. In Section 4.6, we illustrate DiVLA’s adaptability to new embodiments, successfully completing complex tasks that require bimanual coordination. We provide experimental setup and implementation details in the Appendix.
4.1 Experimental Setup.
Implementation details and pretraiend data. The model is pre-trained on the Droid (Khazatsky et al., 2024) dataset. We then finetune our model on evaluation tasks, similar to the setting as $\pi_{0}$ (Black et al., 2024). We use LoRA (Hu et al., 2021) to fine-tune the VLM models. We use 2e-5 as a fixed learning rate to train the model, similar to OpenVLA. The visual encoder and VLM are frozen. We apply LoRA on VLM for fine-tuning.
Data for finetuning. We explore four experimental settings: factory sorting, bin picking, multi-task learning, and table bussing. The first three settings are conducted with the Franka robot, while the table bussing task utilizes the bimanual AgileX robot. Our dataset includes 500 trajectories for the factory sorting task and 580 trajectories for multi-task learning. The bin picking task is designed as a zero-shot task, so no training data was collected for it. For the table bussing task, we gathered 400 trajectories, where objects are randomly placed on the table, often overlapping with each other. During the fine-tuning stage, all data corresponding to the same embodiment are trained together. For instance, the factory sorting and multi-task data are combined for training purposes.